Short Report – QA/QC Meeting June 2018

Meeting Date: 2018-06-14

Participants / Excused

LH, MG, EPL, KF, IF, PM (6) / PL, SB, RM

Talking Points


  1. If you are (Tech-)SRP, please attend QA/QC meetings or tell LH if you can’t, needed for planning of the meetings.
  2. SRP & Tech-SRP: please send short statement about site to LH (max. 2 sentences)
  3. Currently, automatic fluxes are calculated each day for the last 8 days.
  4. Online QA/QC checks started for CH-DAV, CH-FRU, CH-CHA and CH-OE2: Link Next step is to implement this check for CH-LAE, CH-LAS and CH-AWS, please contact LH.
  5. Preview Swiss FluxNet homepage:
  6. You can check the most recent eddy covariance raw data plots in the following folder on our group drive (P), updated daily:


  7. Regarding GIN: RSS feed needed so that all news from GIN are automatically displayed on Swiss FluxNet homepage. DO will contact LH.
  8. Reminder: the new Graphical Meteo Screening Tool by JE:


  1. [DONE] Automatic fluxes will be activated again (goal is April 2018), priority conflicts w/ ICOS. All sites now active.
  2. [IN PROGRESS] SRP QA/QC Weekly checks in online checklists (instead of the server Excel sheet), CH-DAV and CH-FRU started. Next, CH-LAE, CH-LAS and CH-AWS, please contact LH.
  3. [DONE] Activate site-specific RSS feeds. Please contact LH for subscribing to the feed so you can get the newest Swiss FluxNet news for your site. Active subscriptions at the moment: SRPs: CH-CHA, CH-OE2, CH-DAV; T-SRPs: none
  4. [?UPCOMING?] Tech-SRP QA/QC Weekly checks in online checklists. As soon as ICOS allows.
  5. [UPCOMING] On-demand flux calculations on grassland server instead of RDS. As soon as ICOS allows.
  6. [UPCOMING] QV File Reports: checks available files on server. As soon as ICOS allows.

Site Reports


SRP (LH, fluxes): Unfortunate EC data gap mainly in the first half of May (7-17 May, 11 days) during period of increasing CO2 uptake, but otherwise fluxes look good.
Tech-SRP (IB): Data acquisition is working fine, the IRGA needs to be cleaned/calibrated soon. There seems to be a problem with transferring data from the site to our server, same as in CHA, FRU and other sites?


SRP (IF): On  May 30 additional sensors for the IFDewS project have been installed on site by Andreas and Paul. The second harvest of the year took place on June 8 and I (Iris) went to collect biomass samples on the same day. Fluxes look good, still no QCL.  I am planning a reinstrumentation (mainly of the soil sensors) and an extensive soil sampling in mid-July.
Tech-SRP (PL): –


SRP (MG, LH): Since the “Big IRGA Calibration Data Gap of 2018” flux measurements are running well and show consistent CO2 uptake. Data shown in the QuickViews now show the “new” data files that are soon used for ICOS data (previously the non-ICOS files were shown: same data, but different format).
Tech-SRP (PM): Sonic measurements with few day data-gap due to software-testing. Snowframes removed from all chamber collars. CNR4 radiation measurements with sudden change, probably an issue with the heating of the sensor.


SRP (LH): Both EC and meteo data are complete and look good. There is still an issue w/ the variable G_GF1_0x06_2_Avg_Wm-2_Avg, but considering these are replicate measurements the sensor will most likely be dismounted.
Tech-SRP (PL): –


SRP (EPL): The flow from the LI-7200 was obstructed from May 31st to June 7th, the fluxes during that period are therefore unusable (we need a quality flag based on the flow). This year had one of the earliest season start at the end of April at CH-LAE, the forester said this is one of the earliest times the leaves came out in his career of about 30 years.
Tech-SRP (TB): There was a problem with the flow of the LI7200. Flow was zero. After checking the system at the site we noticed that we had the device filled with water and probably therefore the flow modul broke down. Replaced the flow modul. We also had a site visit with Peter Waldner from WSL to discuss the transition of the ownership of the tower from EMPA to WSL.


Tech-SRP (TB): We installed the Laserspectrometer for the Innofarm project. I removed the windvane since it is broken. We are still waiting for replacement from Campbell. Checked out SWC sensor –> seems ok. Ethernet switch still needs a replacement.


You are here: Short Report - QA/QC Meeting June 2018


Last Updated on 14 Feb 2024 12:15

Leave a Reply

You must be logged in to post a comment.